# Post-detection inference for sequential changepoint localization

### v3

@@ C.2 Another concrete example: Poisson rate change problem
In this setting, we have the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (B.13) and the same method can be employed for constructing $1-\gamma$ confidence interval for $\theta_{0}$ as well.
@@
$Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise. }\end{cases}$ (C.14)
@@
Appendix D Experiments for confidence intervals for pre/post change parameters
D.1 Setting II
Table D.1 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (5.1)) for $\theta_{1}$ across 500500500 independent runs with $\alpha=\beta=\eta=0.05$. The data generation and change detection procedure remain the same as in the experiment described in Setting II of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of 0.850.850.85, indicating that our method is reliable but conservative.
Table D.1: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change. The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column of the table. (Setting II)
@@ D.1 Setting II
100 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.98 0.98 1.37 [0.76, 2.13]
100 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.96 1.23 [0.90, 2.16]
500 [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.99 0.93 1.48 [0.75, 2.23]
500 [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.91 1.31 [0.90, 2.21]
D.2 Setting III
Table D.2 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), which are defined in (5.2) and (5.1)), for $\theta_{0}$ and $\theta_{1}$ across 500500500 independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$. he data generation and change detection procedure remain the same as in the experiment described in Setting III of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of 0.800.800.80.
Table D.2: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.850.85(Setting III): unknown pre- and post-change. The true (unknown) values of the parameters are $\theta_{0}=0$ and $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns of the table.
T
$\Theta_{0}$
$\Theta_{1}$
@@ D.2 Setting III
100 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 1.00 0.99 1.00 0.99 0.55 1.19 [-0.32, 0.23] [0.76, 1.95]
100 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.98 0.99 0.98 0.98 0.61 1.16 [-0.51, 0.10] [0.90, 2.06]
500 (,0.25]($\infty$,0.25]( ∞ , 0.25 ] [0.75,)[0.75$\infty$)[ 0.75 , ∞ ) 0.99 0.99 0.93 0.93 0.21 1.22 [-0.10, 0.11] [0.75, 1.97]
500 (,0.1]($\infty$,0.1]( ∞ , 0.1 ] [0.9,)[0.9$\infty$)[ 0.9 , ∞ ) 0.99 0.99 0.94 0.94 0.19 1.24 [-0.09, 0.10] [0.90, 2.14]
D.3 Comparison with [36]
Next, we compare our confidence interval for the post-change parameter with that proposed by [39] and discussed in [36]. We implement Wu’s interval as discussed in Chapter 4.3 of [36] (or Section of [39]), replacing $\theta$ by $\hat{\theta}$ in the bias and variance terms, which is given by $\hat{\theta}-\frac{4}{5\sqrt{d\hat{\theta}(\tau-\hat{\nu})}}\pm z_{\alpha/2}\frac{\sqrt{1-{13}/(16d\hat{\theta})}}{\sqrt{\tau-\hat{\nu}}}$, where $\hat{\theta}=\frac{T_{\tau}}{\tau-\hat{\nu}}$, $\tau=\tau^{\prime}_{\text{cusum}}$, and $z_{\alpha^{\prime}/2}$ is the $\alpha^{\prime}/2$ quantile of the standard normal distribution. We use CUSUM [18] (the same change detector as in the last experiment of Section 6), with $d=10$ for both methods. For our method, we assume the post-change space to be $\Theta_{1}=[0,\infty)$ with $\alpha=\beta=\eta=0.05$. We report the average size and the conditional and unconditional coverage rates of both methods across 500500500 independent runs in Table D.3, with a target conditional coverage of $\alpha^{\prime}=0.85$. The results show that the existing method [36], which is based on asymptotic approximations, undercovers $\theta_{1}$ (both conditionally and marginally). At the same time, our finite-sample confidence intervals are conservative — achieving higher-than-nominal (conditional) coverage at the expense of larger (conditional) length.
Table D.3: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage 0.850.850.85) with Wu [36] (Setting II)
@@ D.3 Comparison with [36]
${N}(-0.25,1)$ ${N}(0.25,1)$ 100 1.46 0.47 0.99 0.77 0.98 0.74
${N}(-0.3,1)$ ${N}(0.3,1)$ 100 1.43 0.51 0.98 0.76 0.98 0.75
${N}(-0.25,1)$ ${N}(0.25,1)$ 500 1.48 0.49 0.99 0.78 0.91 0.64
${N}(-0.3,1)$ ${N}(0.3,1)$ 500 1.49 0.51 0.99 0.79 0.91 0.71
Appendix E Additional experiments for composite pre/post change
@@
$Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})=\begin{cases}\theta_{i}^{(0)}+\epsilon_{n}^{j},\text{ if }n<t\\ \theta_{k}^{(1)}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (E.1)
@@
$\mathcal{C}=\left\{t\in\{1,\cdots,\tau\}:M_{t}\leq\max_{i\in[I],k\in[K]}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{j=1}^{B})\right\}$ (E.2)
@@
As shown in Table E.1, the confidence sets (E.2) have lengths and coverages that are comparable to those from the earlier method reported in Table 6.4 and Table 6.6. However, this grid-based approach incurs significantly higher computational cost. Therefore, for practical purposes, we recommend using the earlier, more computationally efficient method.
Table E.1: Composite pre/post change
@@
Appendix F Simpler confidence sets for detection algorithms controlling PFA
@@
$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha}\right\},$ (F.1)
@@
Theorem F.1.
Let $\mathcal{C}$ be the confidence set as defined in (F.1), for some $\alpha\in(0,1)$. If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, then for any $F_{0}\in\mathcal{P}_{0}$ and $F_{1}\in\mathcal{P}_{1}$,
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha-\delta\leavevmode\nobreak\ \text{, and }\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{1-\delta},$
@@
A similar simplification can be applied to the adaptive confidence sets introduced in Section 4 for the parametric setting. By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 4.1, 4.4 and 4.5, we obtain
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}\}_{j=1}^{B})\right\},$ (F.2)
@@
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}_{\tau-t+1}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})\right\},$ (F.3)
@@
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (F.4)
@@
For detection algorithms with PFA $\leq\delta,\leavevmode\nobreak\ \delta\in(0,1)$, these simpler forms retain the coverage guarantees — both conditional and unconditional — similar to what we have shown in the theorem above.
@@
Appendix G Some additional experimental results
G.1 Experiments with dependent pre- and post-change data
@@
$\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(j:n)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(j:n)}\geq A\right\},$ (G.1)
@@
After detection, we compute the adaptive confidence sets (4.1) with the likelihood ratio-based test statistic (3.2) and point estimate as in (3.1). Note that our adaptive method (4.1) is the only one we know that is valid for dependent data. Table G.1 contains the results averaged across 500500500 runs.
Table G.1: Known pre- and post-change with dependency (Setting I)
@@
G.2 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
G.2.1 Known pre- and post-change (Setting A, PFA 0.001$\leq$ 0.001≤ 0.001)
@@ G.2 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (G.2)
[…]
Table G.2: Known pre- and post-change distributions (Setting A, PFA 0.001$\leq$ 0.001≤ 0.001)
@@
[…]
G.2.2 Known pre- and unknown post-change (Setting B, PFA 0.001$\leq$ 0.001≤ 0.001)
@@
$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (G.3)
with thresholds at $A=1000$. The weights are chosen just as described in Table 6.2. Similarly, the PFA of this detector is at most $1/A$, i.e., 0.0010.0010.001 here. After detection, we construct our universal confidence sets (F.1) for $T$, with $\alpha=0.075$. The test statistic and the point estimate are the same as mentioned in Setting II of Section 6. We report the average size and the conditional and unconditional coverage rates across 500500500 independent runs in Table G.3. We also report the average absolute deviation of $\hat{T}$ in (B.2) and delay (given no false detection) in the table.
Table G.3: Known pre-change but unknown post-change (Setting B, PFA 0.001$\leq$ 0.001≤ 0.001)
@@
G.2.3 Composite pre- and post-change (Setting C, PFA 0.001$\leq$ 0.001≤ 0.001)
@@
$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (G.4)
[…]
Table G.4: Composite pre- and post-change (Setting C, PFA 0.001$\leq$ 0.001≤ 0.001)
@@
G.2.4 Huber-robust (nonparametric) pre- and post-change (Setting D, PFA 0.001$\leq$ 0.001≤ 0.001)
We perform experiments for the robust Gaussian mean-change problem under the total variation contamination model, with the true changepoint at $T=100$ and $T=500$. The pre- and post-change classes are total variation balls of radius $\epsilon$ around $N(\mu_{0},1)$ and $N(\mu_{1},1)$, respectively.
$\mathcal{P}_{i}=\{F\in\mathcal{M}:d_{TV}(F,N(\mu_{i},1))\leq\epsilon\};\leavevmode\nobreak\ \leavevmode\nobreak\ i=0,1.$
In our experiment, $\mu_{0}=0,\mu_{1}=1$ and $\epsilon=0.01,0.001$. The true (unknown) pre- and post-change distributions are $(1-\epsilon)\times N(0,1)+\epsilon\times\text{Cauchy}(-1,10)$ and $(1-\epsilon)\times N(1,1)+\epsilon\times\text{Cauchy}(-1,10)$ respectively. We use the Huber-robust tests by [23] as change-detectors:
$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}E_{i}\geq A\right\},$ (G.5)
[…]
After detection, we construct our universal confidence sets (F.1) for $T$, with $R_{n}^{(t)}=\prod_{i=t}^{n}E^{\prime}_{i}$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}E_{i}$, for $n<t$ and $\alpha=0.05$, where $E^{\prime}_{i}=\frac{\pi^{\prime}(X_{i})}{\mathbb{E}_{F_{1}}\pi^{\prime}(X_{i})+(c^{\prime\prime}-c^{\prime})\epsilon},$ $\pi^{\prime}(x)=\min\{\max\{\frac{f_{0}(x)}{f_{1}(x)},c^{\prime}\},c^{\prime\prime}\}$ and $E_{i}$ is as defined before. We report the average size and coverage rates across 200200200 independent runs in Table G.5. We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection). The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
Table G.5: Huber-robust (nonparametric) pre- and post-change (Setting D, PFA 0.001$\leq$ 0.001≤ 0.001)
T $\epsilon$
Conditional
coverage
Marginal
coverage
Size
$|\hat{T}-T|$
Delay
100 0.01 0.97 0.97 16.04 3.23 119.38
100 0.001 0.97 0.97 14.57 3.18 113.06
500 0.01 0.96 0.96 16.18 3.87 536.58
500 0.001 0.97 0.97 15.13 3.25 531.44
G.2.5 Comparison of experimental results with ARL and PFA controlling detection algorithms

### v4

@@ D.2 Another concrete example: Poisson rate change problem
In this setting, we have the $1-\beta$ confidence sequence for $\theta_{1}$, as defined in (S13) and the same method can be employed for constructing $1-\gamma$ confidence interval for $\theta_{0}$ as well.
@@
$Y_{n}^{j}(\theta,\theta^{\prime}):=\begin{cases}\sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta/\lambda_{t}),\text{ if }n<t\\ \sum_{i=1}^{X_{n}^{j}}\mathds{1}(U_{n,i}^{j}<\theta^{\prime}/\lambda_{t}),\text{ otherwise. }\end{cases}$ (S14)
@@
Appendix E Experiments for confidence intervals for pre/post change parameters
E.1 Setting II
Table E.1 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries) (defined in (5.1)) for $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta=0.05$. The data generation and change detection procedure remain the same as in the experiment described in Setting II of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.85$, indicating that our method is reliable but conservative.
Table E.1: Confidence interval for $\theta_{1}$: known pre-change but unknown post-change. The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column of the table. (Setting II)
@@ E.1 Setting II
100 $0.75,\infty)$ 0.98 0.98 1.37 [0.76, 2.13]
100 $0.9,\infty)$ 0.98 0.96 1.23 [0.90, 2.16]
500 $0.75,\infty)$ 0.99 0.93 1.48 [0.75, 2.23]
500 $0.9,\infty)$ 0.98 0.91 1.31 [0.90, 2.21]
E.2 Setting III
Table E.2 contains conditional and marginal (empirical) coverages, (conditional) average of the sizes and of the confidence intervals (i.e., the upper and lower boundaries), and (5.1)), for $\theta_{0}$ and $\theta_{1}$ across $500$ independent runs with $\alpha=\beta=\eta_{0}=\eta_{1}=0.05$. he data generation and change detection procedure remain the same as in the experiment described in Setting III of Section 6. Across all settings, the observed conditional coverage consistently exceeds the theoretical lower bound of $0.80$.
Table E.2: Confidence intervals for $\theta_{0},\theta_{1}$ with target conditional coverages 0.850.85(Setting III): unknown pre- and post-change. The true (unknown) values of the parameters are $\theta_{0}=0$ and $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns of the table.
T $\Theta_{0}$ $\Theta_{1}$
@@ E.2 Setting III
100 $(\infty,0.25]$ $0.75,\infty)$ 1.00 0.99 1.00 0.99 0.55 1.19 [-0.32, 0.23] [0.76, 1.95]
100 $(\infty,0.1]$ $0.9,\infty)$ 0.98 0.99 0.98 0.98 0.61 1.16 [-0.51, 0.10] [0.90, 2.06]
500 $(\infty,0.25]$ $0.75,\infty)$ 0.99 0.99 0.93 0.93 0.21 1.22 [-0.10, 0.11] [0.75, 1.97]
500 $(\infty,0.1]$ $\infty) 0.99 0$.99 0.94 0.94 0.19 1.24 [-0.09, 0.10] [0.90, 2.14]
E.3 Comparison with [39]
Next, we compare our confidence interval for the post-change parameter with that proposed by [42] and discussed in [39]. We implement Wu’s interval as discussed in Chapter 4.3 of [39] (or Section of [42]), replacing $\theta$ by $\hat{\theta}$ in the bias and variance terms, which is given by $\hat{\theta}-\frac{4}{5\sqrt{d\hat{\theta}(\tau-\hat{\nu})}}\pm z_{\alpha/2}\frac{\sqrt{1-{13}/(16d\hat{\theta})}}{\sqrt{\tau-\hat{\nu}}}$, where $\hat{\theta}=\frac{T_{\tau}}{\tau-\hat{\nu}}$, $\tau=\tau^{\prime}_{\text{cusum}}$, and $z_{\alpha^{\prime}/2}$ is the $\alpha^{\prime}/2$ quantile of the standard normal distribution. We use CUSUM [19] (the same change detector as in the last experiment of Section 6), with $d=10$ for both methods. For our method, we assume the post-change space to be $\Theta_{1}=[0,\infty)$ with $\alpha=\beta=\eta=0.05$. We report the average size and the conditional and unconditional coverage rates of both methods across $500$ independent runs in Table E.3, with a target conditional coverage of $\alpha^{\prime}=0.85$. The results show that the existing method [39], which is based on asymptotic approximations, undercovers $\theta_{1}$ (both conditionally and marginally). At the same time, our finite-sample confidence intervals are conservative — achieving higher-than-nominal (conditional) coverage at the expense of larger (conditional) length.
Table E.3: Comparison of the confidence sets for $\theta_{1}$ (target conditional coverage $0.85$) with Wu [39] (Setting II)
@@ E.3 Comparison with [39]
N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 100 1.46 0.47 0.99 0.77 0.98 0.74
N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 100 1.43 0.51 0.98 0.76 0.98 0.75
N(0.25,1){N}(-0.25,1) N(0.25,1){N}(0.25,1) 500 1.48 0.49 0.99 0.78 0.91 0.64
N(0.3,1){N}(-0.3,1) N(0.3,1){N}(0.3,1) 500 1.49 0.51 0.99 0.79 0.91 0.71
Appendix F Additional experiments for composite pre/post change
@@
$Y_{n}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})=\begin{cases}\theta_{i}^{(0)}+\epsilon_{n}^{j},\text{ if }n<t\\ \theta_{k}^{(1)}+\epsilon_{n}^{j},\text{ otherwise. }\end{cases}$ (S1)
@@
$\mathcal{C}=\left\{t\in\{1,\cdots,\tau\}:M_{t}\leq\max_{i\in[I],k\in[K]}\operatorname{Quantile}(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta_{i}^{(0)},\theta_{k}^{(1)})\}_{j=1}^{B})\right\}$ (S2)
@@
As shown in Table F.1, the confidence sets (S2) have lengths and coverages that are comparable to those from the earlier method reported in Table 6.3 and Table 6.5. However, this grid-based approach incurs significantly higher computational cost. Therefore, for practical purposes, we recommend using the earlier, more computationally efficient method.
Table F.1: Composite pre/post change
@@
Appendix G Simpler confidence sets for detection algorithms controlling PFA
@@
$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha}\right\},$ (S1)
@@
Theorem G.1.
Let $\mathcal{C}$ be the confidence set as defined in (S1), for some $\alpha\in(0,1)$. If $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$, then for any $F_{0}\in\mathcal{P}_{0}$ and $F_{1}\in\mathcal{P}_{1}$,
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha-\delta\penalty 10000\ \text{, and }\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{1-\delta},$
@@
A similar simplification can be applied to the adaptive confidence sets introduced in Section 4 for the parametric setting. By replacing the $\alpha r_{t}^{*}$ term with $\alpha$ in Equations 4.1, 4.3 and 4.4, we obtain
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}\}_{j=1}^{B})\right\},$ (S2)
@@
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}_{\tau-t+1}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})\right\},$ (S3)
@@
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (S4)
@@
For detection algorithms with PFA $\leq\delta,\penalty 10000\ \delta\in(0,1)$, these simpler forms retain the coverage guarantees — both conditional and unconditional — similar to what we have shown in the theorem above.
@@
Appendix H Some additional experimental results
H.1 Experiments varying $T$ and ARL
We have seen in deriving the asymptotic length bound that it increases with the ratio $T/A_{T}$. Now, we experimentally validate the fact by conducting the same experiment as in Setting I (known pre- and post-change) in Section 6 with varying $T$ thresholds $A$ of the CUSUM detector (larger threshold means large ARL). Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 3.5 are reported in Table H.1. The results show a clear pattern: as the ratio $T/A$ increases, the resulting confidence sets become wider, reflecting growing conservativeness.
Table H.1: Setting I: Known pre- and post-change. Average (of $500$ independent runs) size and coverage of the confidence sets using Equation 3.5 with $F_{0}=N(0,1),F_{1}=N(1,1),N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ (3.1) and average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
$T$ A
Conditional
coverage
Marginal
coverage
Size
$|\hat{T}-T|$
Delay
50 100 0.98 0.91 17.28 2.58 9.33
1000 100 0.99 0.21 20.88 2.45 9.14
2000 100 1 0.07 27.02 2.60 8.35
500 1000 0.98 0.97 15.77 2.62 13.22
10000 1000 1 0.27 24.29 2.36 13.25
20000 1000 1 0.07 30.16 2.29 13.22
H.2 Experiments with dependent pre- and post-change data
@@
$\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\left(\frac{p_{1}}{p_{0}}\right)^{n_{10}(j:n)}\left(\frac{1-p_{1}}{1-p_{0}}\right)^{n_{11}(j:n)}\geq A\right\},$ (S1)
@@
After detection, we compute the adaptive confidence sets (4.1) with the likelihood ratio-based test statistic (3.2) and point estimate as in (3.1). Note that our adaptive method (4.1) is the only one we know that is valid for dependent data. Table H.2 contains the results averaged across $500$ runs.
Table H.2: Known pre- and post-change with dependency (Setting I)
@@
H.3 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
H.3.1 Known pre- and post-change (Setting A, PFA $\leq 0.001$)
@@ H.3 Experiments with detection algorithms that control PFA for Gaussian mean change scenario
$\tau=\inf\left\{t\in\mathbb{N}:\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ (S2)
with thresholds at $A=1000$. Note that $\prod_{i=1}^{t}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$ is a non-negative martingale under $T=\infty$ and hence, by Ville’s inequality [35], the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we construct both the proposed confidence sets (universal (S1) and adaptive (S2)), with $B=100$, $\alpha=0.05$. We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.3. We also report the average absolute deviation of $\hat{T}$ (defined in (3.1)) and delay (given no false detection) in the table.
Table H.3: Known pre- and post-change distributions (Setting A, PFA $\leq 0.001$)
@@
500 $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 18.89 16.82 3.18 417.44
H.3.2 Known pre- and unknown post-change (Setting B, PFA $\leq 0.001$)
@@
$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (S3)
with thresholds at $A=1000$. The weights are chosen just as described in Table 6.2. Similarly, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. After detection, we construct our universal confidence sets (S1) for $T$, with $\alpha=0.075$. The test statistic and the point estimate are the same as mentioned in Setting II of Section 6. We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.4. We also report the average absolute deviation of $\hat{T}$ in (S2) and delay (given no false detection) in the table.
Table H.4: Known pre-change but unknown post-change (Setting B, PFA $\leq 0.001$)
@@
H.3.3 Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
@@
$\tau=\inf\left\{t\in\mathbb{N}:\int_{\theta\in\Theta_{1}}\prod_{i=1}^{t}\frac{f_{\theta}(X_{i})}{f^{*}_{\theta}(X_{i})}dW(\theta)\geq A\right\},$ (S4)
with $f^{*}_{\theta}$ being the RIPr of $f_{\theta}$ on the pre-change class $\mathcal{P}_{0}$ and with thresholds at $A=1000$. This is a stopping time of an “e-process” [20] and hence, the PFA of this detector is at most $1/A$, i.e., $0.001$ here. The weights are chosen just as described in Table 6.3. After detection, we construct our universal confidence sets (S1) for $T$, with $\alpha=0.1$. The test statistic and the point estimate are the same as mentioned in Setting III of Section 6. We report the average size and coverage rates across $500$ independent runs in Table H.5. We also report the average absolute deviation of $\hat{T}$ in (S2) and delay (given no false detection). The sizes and coverages are reasonable, despite the methods not knowing the data-generating distribution exactly.
Table H.5: Composite pre- and post-change (Setting C, PFA $\leq 0.001$)
@@
H.3.4 Comparison of experimental results with ARL and PFA controlling detection algorithms